# Recent Advances in Discrete Speech Tokens: A Review

### v2

@@ IV-E Variable Frame Rate Tokens and Unit Discovery
Information in speech is not uniformly distributed along the time axis [163]. In segments such as silence or long vowels, information density is low, whereas in segments with explosive consonants, speech events occur much more frequently. This inherent non-uniformity suggests that it might be more natural to allocate more tokenized bits to regions with dense information and higher variance, and fewer bits to regions with less uncertainty. This kind of discrete speech tokens is referred to as variable frame rate (VFR) tokens in this review. Note that while multi-resolution and variable-bitrate tokens have been introduced previously, the concept of VFR is still distinct. In multi-resolution tokens [23, 74], each quantizer operates at a fixed frame rate. In variable-bitrate tokens [24], the frame rate remains fixed, while the variability lies in the number of quantizers per frame. Instead, VFR tokens should directly allocate different granularities on the temporal axis.
@@
VFR tokens are closely related to acoustic unit discovery. As speech lacks a natural boundary of phonetic units [8], there are much research efforts to find and locate the underlying acoustic units behind speech utterances in an unsupervised manner [164, 165, 166, 167]. This is particularly of interest for low-resource languages. The discovered units can guide the boundary segmentation of VFR tokens. To this end, VFR tokens are interesting not only because they might reduce the necessary bitrate, but also because they can introduce a strong inductive bias that linguistic knowledge is encoded [163].
@@
A recent direction of VFR tokens is to discover acoustic units from an SSL model. Note that deduplicated tokens and acoustic BPE themselves can be regarded as VFR tokens. Sylber [168] and SyllableLM [45] take similar approaches that first locate acoustic boundaries from existing HuBERT models, and then train another HuBERT student with segment-level pooled targets between boundaries. The final HuBERT embeddings undergo the same segment-level pooling and kmeans clustering procedure to produce tokens at a very low frame rate ($\approx 5$Hz) that align well with syllables.
@@
Boundary prediction can be involved to achieve frame rate variability in the training process, where a specific model predicts frame-level boundaries and is trained together with other network modules. The training techniques of such models include reinforcement learning [169], soft pooling [148], and slowness constraint [163]. However, these approaches are rarely adopted in the context of discrete speech tokens today, and there have barely been a VFR acoustic token till now.
IV-F Speech Token Vocoders
Acoustic tokens are designed naturally with a decoder that outputs waveforms or spectrograms given tokens, but semantic tokens are not. A necessary component for building a discrete token-based speech generation system with semantic tokens is the speech resynthesis model, or speech token vocoders. Unlike traditional spectrogram-based vocoders [170], these vocoders receive discrete speech tokens as an input and reconstruct speech signals.
Polyak et al. [143] first explores speech resynthesis from discrete speech units by a HifiGAN [170] augmented with discretized pitch units and speaker embedding inputs. The vec2wav vocoder in VQTTS [129] improves this vocoder by a Conformer [93] frontend module before HifiGAN generator. Later, CTX-vec2wav [111] proposes a position-agnostic cross-attention mechanism that effectively integrates timbre information from surrounding acoustic contexts without the need for pretrained speaker embeddings. This makes it more timbre-controllable and suitable for zero-shot TTS and VC [112]. Upon it, vec2wav 2.0 [171] further advances the timbre controllability by SSL timbre features and adaptive activations, demonstrating a strong VC performance.
It is also feasible to apply diffusion or flow matching algorithms in token vocoders [172, 173, 41]. There, the discrete tokens are treated as a condition for diffusion or flow matching to generate mel-spectrograms, and further converted to waveform by a pretrained mel vocoder. Compared to training a token-to-wav vocoder in an end-to-end fashion, training a token-to-mel model is more convenient and does not need adversarial training. To better control timbre, a mask strategy is introduced into the training process where the model only computes loss on the un-masked part of spectrograms [41]. During inference, spectrogram from speaker prompt conditions the generative process, which can be regarded as a form of “in-context learning”. However, this requires tokens to be extracted from reference prompts before synthesis. Also, inference efficiency may be compromised for better generation quality with multiple inference steps, and this method is only validated on massive amount of data currently.
V Analysis of Discrete Speech Tokens
V-A Metrics and Existing Benchmarks
@@
  • Signal-level reconstruction metrics: For reconstruction evaluations, signal-level metrics like PESQ, STOI, mel distance, GPE, etc. are often used.
@@
  • Perceptual reconstruction metrics: Apart from signal-level metrics, there can also be perceptual evaluations of reconstruction performance. This includes intelligibility (often measured by WER, CER, or phone error rates), speaker similarity, subjective or proxy MOS tests, etc.
@@
  • Performance on downstream tasks: Probing tasks can be used to measure the preservation or prominence of certain information in tokens, like ASR, ASV, emotion recognition, and spoken language modeling [167]. Note that this is different from perceptual reconstruction metrics since it operates directly on tokens. Performance in generative tasks like TTS and VC can also be evaluated.
@@
  • Semantic/phonetic relevance: If the tokens are expected to align with texts (e.g. for semantic tokens and semantic-distilled acoustic tokens), metrics like phone discrminability [167], phone purity, and phone-normalized mutual information [33] can be computed.
@@
  • Invariance and robustness: If the tokens are expected to be invariant to perturbations, unit edit distance [39] can be considered as a measurement.
@@
There are several existing benchmarks on discrete speech tokens. Codec-SUPERB [^5] https://codecsuperb.com/ [174] evaluates both signal-level reconstruction metrics and downstream performances of acoustic tokens. ESPnet-Codec [175] integrates multiple codecs into a unified training and evaluation framework VERSA [176] and extends evaluation to some generative tasks such as TTS and SVS. DASB [^6] https://github.com/speechbrain/benchmarks/tree/main/benchmarks/DASB [144] performs more downstream probing tasks, and includes generative tasks as well as semantic tokens. STAB [177] takes a different perspective that measures the invariance, robustness, compressibility, and vocabulary utilization of speech tokens. This emphasizes the application in spoken language models instead of reconstruction.
V-B Existing Analyses
@@
There are several theoretical or experimental analyses of the advantages of discrete speech tokens. Nguyen et al. [178] demonstrates by an encoder-only language model that semantic SSL tokens are favorable for spoken language modeling, due to their removal of linguistically irrelevant information. Sicherman et al. [115] supports this claim by showing that semantic units have a strong correlation with phonemes, but a weaker correlation with speakers. Abdullah et al. [179] refines the correlation between semantic SSL tokens and linguistic content to the “sub-phonemic” level instead of high-level phonemes due to contextual variability. Chang et al. [128] explores the use of WavLM tokens for end-to-end ASR together with deduplication and BPE. Although these tokens underperform continuous SSL features, they still show competitive performance. Similar findings are reported on contextual ASR [180], multilingual ASR [181], end-to-end speech translation, understanding [156], and more LLM-based semantic-related tasks with discrete units as inputs [182]. Expresso [183] evaluates the resynthesis quality of HuBERT and EnCodec tokens on an expressive dataset, finding that HuBERT struggles to preserve source speech expressivity while EnCodec performs better.
@@
The downside of speech tokens is also researched. Yeh et al. [116] suggests that VQ on HuBERT embeddings cannot achieve perfect disentanglement of speaker and phonetic content. EMO-Codec [184] shows that codec reconstruction still sometimes degrades the emotion information. O’Reilly et al. [185] shows that neural audio codecs often lack stability after repeated encoding and decoding, i.e. not idempotent.
@@
Therefore, the reconstruction quality of acoustic tokens and the performance on discriminative downstream tasks of both acoustic and semantic tokens have been well benchmarked. However, the reconstruction performance of semantic tokens still requires a more thorough comparison. Hence, we adopt a reconstruction approach to compare different types of tokens. Specifically, we use a timbre-controllable speech token vocoder to resynthesize semantic tokens into waveforms and measure the preservation of content, prosody, speaker identity, and acoustic details, respectively. The details and results of these experiments will be elaborated in the rest of this chapter.
V-C Reconstruction Analysis
@@
TABLE III: Reconstruction and voice conversion comparisons of tokens in different categories. For all semantic tokens, we train a specific CTX-vec2wavα as a vocoder. All vocoders are trained on LibriTTS [186], and evaluations are done on LibriTTS testset-B [111]. “L” means a certain layer in the SSL Transformer block, and “km” means manual k-means clustering. Settings in parentheses denote model versions.
@@
Continuous Baselines Ground Truth Recording - 1.16 0.00 4.50 1.000 1.000 - - -
Mel + BigVGAN (100 band) - 1.18 0.88 4.30 0.995 0.997 - - -
Acoustic Token (Reconstruction Only) EnCodec ($Q=8$) 6.00 1.53 1.33 2.83 0.946 0.979 - - -
DAC (24kHz, $Q=8$) 6.00 1.34 0.93 3.52 0.958 0.982 - - -
TiCodec ($Q=2$) 1.50 3.31 1.51 2.00 0.898 0.905 2.62 0.642 0.886
SNAC (24kHz) 0.98 2.25 1.48 2.23 0.914 0.952 - - -
WavTokenizer (Small, $F=75$Hz) 0.90 2.45 1.63 2.47 0.925 0.960 - - -
Stable-Codec ($Q=12$) 0.70 4.94 1.73 2.16 0.917 0.889 - - -
Acoustic Token (Semantic Distillation) SpeechTokenizer 4.00 1.47 1.20 2.60 0.930 0.972 - - -
X-Codec (HuBERT LibriSpeech) 4.00 1.27 1.49 2.82 0.905 0.971 - - -
Mimi 1.10 2.44 1.68 2.27 0.917 0.938 - - -
LLM-Codec 0.85 6.25 1.86 1.82 0.879 0.919 - - -
SemantiCodec ($F$=25Hz, $V$=$2^{13}$+$2^{15}$) 0.70 3.44 2.28 1.75 0.866 0.930 - - -
Acoustic Token (Disentanglement) FACodec (with detail codes) 4.80 1.37 1.02 2.91 0.954 0.971 1.57 0.773 0.583
LSCodec ($F=50$Hz) 0.45 3.33 2.42 1.77 0.688 0.954 4.04 0.852 0.697
Semantic Token (SSL) vq-wav2vec (k-means) 1.80 2.81 2.73 1.49 0.795 0.940 3.27 0.857 0.718
wav2vec 2.0 (Large, inner quantizer) 0.90 3.24 2.92 1.52 0.680 0.947 4.40 0.814 0.759
wav2vec 2.0 (Large, L14+km2048) 0.55 2.51 9.57 1.20 0.630 0.933 2.81 0.880 0.492
HuBERT (Large, L24+km2048) 0.55 1.86 15.65 1.17 0.625 0.934 1.97 0.876 0.375
WavLM (Large, L24+km2048) 0.55 1.67 17.94 1.16 0.621 0.934 1.92 0.872 0.374
Semantic Token (SSL+Perturb Invariance) ContentVec (L12+km2048) 0.55 2.09 18.88 1.15 0.613 0.921 2.21 0.869 0.348
Semantic Token (Supervised) $\mathcal{S}^{3}$ Tokenizer ($F=50$Hz) 0.60 2.12 4.25 1.37 0.673 0.944 2.52 0.868 0.687
To enable a fair comparison between acoustic and semantic tokens from a reconstruction perspective, we train a CTX-vec2wavα vocoder [29] for different semantic tokens on LibriTTS [186]. This vocoder supplements the insufficient speaker timbre information in semantic tokens using continuous WavLM features extracted from the reference prompts. This approach enables semantic tokens to perform voice conversion (VC) by switching reference prompts conveniently. The training details follow [111]. We compute several metrics for reconstruction ability:
@@
  • SECS (speaker embedding cosine similarity) computes the cosine similarity of speaker embeddings outputted from a speaker verification model [^8] https://github.com/resemble-ai/Resemblyzer.
We use LibriTTS testset-B  [111] as the test set for evaluations. It contains 500 utterances from unseen speakers that sum up to about 1 hour. We use the original utterance to provide timbre information when necessary, i.e. for TiCodec, FACodec, LSCodec and all semantic tokens [^9] Note that there is potential of information leakage in this situation. An alternative strategy is to use a reference prompt from the same speaker to provide timbre information, then compute similarity between the generated utterance and the reference. However, this is rather “same-speaker conversion” instead of reconstruction, and the SECS values are not comparable to ordinary acoustic tokens that don’t require additional speaker inputs.. All evaluation metrics are computed on 16kHz waveform for a fair comparison, and reconstructed waveforms with higher sampling rates are downsampled before evaluation.
@@
The results are shown in Table III. It is evident that acoustic tokens designed only for reconstruction can achieve decent speech quality, but still far from the state-of-the-art spectrogram-based vocoders because of higher compression rates. Retaining good speech intelligibility (i.e. low WER) becomes particularly challenging when the frame rate is low. Acoustic tokens with semantic distillation can also achieve strong reconstruction quality. Explicitly disentangled acoustic tokens may sacrifice some reconstruction performance metrics when the bitrate is extremely low. Semantic tokens generally struggle to achieve the same level of acoustic reconstruction as acoustic tokens, as evidenced by lower GPE, PESQ, and STOI scores. Notably, most semantic tokens included exhibit significant information loss in prosody as reflected by their GPE scores. However, their WER scores remain comparable to acoustic tokens, despite having much lower bitrates. This highlights the property that semantic tokens primarily retain content-related information rather than acoustic details.
V-D Voice Conversion Analysis

### v3

@@ V-B Variable Frame Rate Tokens and Unit Discovery
Information in speech is not uniformly distributed along the time axis [dieleman2021variable]. In segments such as silence or long vowels, information density is low, whereas in segments with explosive consonants, speech events occur much more frequently. This inherent non-uniformity suggests that it might be more natural to allocate more tokenized bits to regions with dense information and higher variance, and fewer bits to regions with less uncertainty. This kind of discrete speech tokens is referred to as variable frame rate (VFR) tokens in this review. Note that while multi-resolution and variable-bitrate tokens have been introduced previously, the concept of VFR is still distinct. In multi-resolution tokens [Siuzdak_SNAC_Multi-Scale_Neural_2024, guo2024speaking], each quantizer operates at a fixed frame rate. In variable-bitrate tokens [chae2024variable], the frame rate remains fixed, while the variability lies in the number of quantizers per frame. Instead, VFR tokens should directly allocate different granularities on the temporal axis.
@@
VFR tokens are closely related to acoustic unit discovery. As speech lacks a natural boundary of phonetic units [mohamed2022self], there are much research efforts to find and locate the underlying acoustic units behind speech utterances in an unsupervised manner [eloff19_interspeech, dunbar20_interspeech, niekerk20b_interspeech, nguyen2020zero]. This is particularly of interest for low-resource languages. The discovered units can guide the boundary segmentation of VFR tokens. To this end, VFR tokens are interesting not only because they might reduce the necessary bitrate, but also because they can introduce a strong inductive bias that linguistic knowledge is encoded [dieleman2021variable].
@@
A recent direction of VFR tokens is to discover acoustic units from an SSL model. Sylber [cho2024sylber] and SyllableLM [baade2024syllablelm] take similar approaches that first heuristically locate acoustic boundaries from existing HuBERT models, and then train another HuBERT student with segment-level pooled targets between boundaries. The final HuBERT embeddings undergo the same segment-level pooling and kmeans clustering procedure to produce tokens at a very low frame rate ($\approx 5$Hz). Such tokens are reported to align well with syllables [baade2024syllablelm, cho2024sylber].
@@
Boundary prediction can be involved to achieve frame rate variability in the training process, where a specific model predicts frame-level boundaries and is trained together with other network modules. The training techniques of such models include reinforcement learning [cuervo2022variable], soft pooling [hwang2024removing], and slowness constraint [dieleman2021variable].
For VFR acoustic tokens, heuristic downsampling methods have been explored recently. TFC [zhang25k_interspeech_tfc] incorporates a frame-rate selection strategy based on entropy values, assigning different downsampling rates to segments with varying entropy levels under a shared quantizer. CodecSlime [wang2025codecslime] adopts a dynamic programming-based downsampling strategy driven by pairwise latent embedding distances, and further introduces a specialized training method to adapt fixed frame-rate codec models to dynamic frame rates. Both approaches enable flexible frame-rate allocation within a single model, and are orthogonal to the codec architecture. However, the study of VFR acoustic tokens remains generally insufficient.
Similar to supervised semantic tokens, explicit linguistic boundaries can also be incorporated into VFR design. Karapiperis et al. [sotirios2024investigating] and TASTE [tseng2025taste] are an initial explorations in this direction, where the tokens are aligned with the frame rate of phonemes or text BPE units, resulting in variable temporal spans. Karapiperis et al. [sotirios2024investigating] is an acoustic token that directly performs phoneme-level downsampling before quantization. In contrast, TASTE aggregates information from a whole spoken utterance onto each text BPE unit by cross-attention. The BPE-level aggregated speech embeddings are then quantized and optimized using a TTS-style objective, These works redefine the notion of tokens: rather than encoding spoken information by themselves, they serve as an additional acoustic layer conditioned on explicit text content.
VI Analysis of Discrete Speech Tokens
VI-A Metrics and Existing Benchmarks
@@
  • Signal-level reconstruction metrics: For reconstruction evaluations, signal-level metrics like PESQ [rix2001perceptual], STOI [taal2011algorithm], mel distance, GPE [bagshaw1993], etc. are often used.
@@
  • Perceptual reconstruction metrics: Apart from signal-level metrics, there can also be perceptual evaluations of reconstruction performance. This includes intelligibility (often measured by word, character, or phone error rates), speaker similarity, subjective listening tests, etc.
@@
  • Performance on downstream tasks: Probing tasks can be used to measure the preservation or prominence of certain information in tokens, like ASR, ASV, emotion recognition, and spoken language modeling [nguyen2020zero]. Note that this is different from perceptual reconstruction metrics since it operates directly on tokens. Performance in generative tasks like TTS and VC can also be evaluated.
@@
  • Semantic/phonetic relevance: If the tokens are expected to align with texts (e.g. for semantic tokens and semantic-distilled acoustic tokens), metrics like phone discrminability [nguyen2020zero], phone purity [hsu2021hubert], and phone-normalized mutual information [hsu2021hubert] can be computed.
@@
  • Invariance and robustness: If the tokens are expected to be invariant to perturbations, unit edit distance [gat2023augmentation] can be considered as a measurement.
@@
There are several existing benchmarks on discrete speech tokens. Codec-SUPERB [wu2024codec] evaluates both signal-level reconstruction metrics and downstream performances of acoustic tokens. ESPnet-Codec [shi2024espnet] integrates multiple codecs into a unified training and evaluation framework VERSA [shi2024versa] and extends evaluation to some generative tasks such as TTS and SVS. DASB [mousavi2024dasb] performs more downstream probing tasks, and includes generative tasks as well as semantic tokens. STAB [vashishth2024stab] takes a different perspective that measures the invariance, robustness, compressibility, and vocabulary utilization of speech tokens. This emphasizes the application in spoken language models instead of reconstruction.
VI-B Existing Analyses
@@
There are several theoretical or experimental analyses of the advantages of discrete speech tokens. Nguyen et al. [nguyen2022discrete] demonstrates by an encoder-only language model that semantic SSL tokens are favorable for spoken language modeling, due to their removal of linguistically irrelevant information. Sicherman et al. [sicherman2023analysing] supports this claim by showing that semantic units have a strong correlation with phonemes, but a weaker correlation with speakers. Abdullah et al. [abdullah23_interspeech] refines the correlation between semantic SSL tokens and linguistic content to the “sub-phonemic” level instead of high-level phonemes due to contextual variability. Chang et al. [chang23b_interspeech] explores the use of WavLM tokens for end-to-end ASR together with deduplication and BPE. Although these tokens underperform continuous SSL features, they still show competitive performance. Similar findings are reported on contextual ASR [cui2024exploring_context], multilingual ASR [cui2024exploring], end-to-end speech translation, understanding [chang2024exploring], and more LLM-based semantic-related tasks with discrete units as inputs [wang2024comparative]. Expresso [expresso] evaluates the resynthesis quality of HuBERT and EnCodec tokens on an expressive dataset, finding that HuBERT struggles to preserve source speech expressivity while EnCodec performs better.
@@
The downside of speech tokens is also studied. Yeh et al. [yeh2024estimating] suggests that VQ on HuBERT embeddings cannot achieve perfect disentanglement of speaker and phonetic content. EMO-Codec [ren2024emo] shows that codec reconstruction still sometimes degrades the emotion information. O’Reilly et al. [o2024code] shows that neural audio codecs often lack stability after repeated encoding and decoding, i.e. not idempotent. ERSB [wang2025towards-asru] reveals speech codecs still struggle under complex acoustic environments, in terms of both reconstruction and downstream task consistency.
@@
Therefore, the reconstruction quality of acoustic tokens and the performance on discriminative downstream tasks of both acoustic and semantic tokens have been benchmarked. However, the reconstruction performance of semantic tokens still requires a more thorough comparison. Hence, we adopt a reconstruction approach to compare different types of tokens. Specifically, we use a timbre-controllable speech token vocoder to resynthesize semantic tokens into waveforms and measure the preservation of content, prosody, speaker identity, and acoustic details, respectively. We also follow DASB [mousavi2024dasb] to conduct semantic modeling probing tasks on various tokens. The detailed setups of these experiments can be found in the appendix.
VI-C Reconstruction Analysis
@@
TABLE I: Reconstruction, voice conversion, and downstream semantic modeling comparisons of tokens in different categories. For reconstruction and voice conversion, we train a specific CTX-vec2wavα vocoder [guo2024lscodec] on LibriTTS [libritts] for all semantic tokens. Settings in parentheses denote model versions. Semantic modeling tasks, ASR and IC (intent classification), follow the DASB [mousavi2024dasb] setup which uses token indices to train a simple LSTM network. “L” means a certain layer in the SSL Transformer block, “km” means manual k-means clustering, and “NC” is short for “not converged”. Please refer to the appendix for evaluation details.
Token Type
Token
Model (version)
@@
Continuous Baselines Ground Truth Recording - 1.16 0.00 4.50 1.000 - - - - -
Mel + BigVGAN (100 band) - 1.18 0.88 4.30 0.995 - - - 17.4 50.1
Acoustic
(General-Purpose)
EnCodec ($Q=8$) 6.00 1.53 1.33 2.83 0.946 - - - 19.4 34.8
DAC (24kHz, $Q=8$) 6.00 1.34 0.93 3.52 0.958 - - - 26.1 18.3
TiCodec ($Q=2$) 1.50 3.31 1.51 2.00 0.898 2.62 0.642 0.886 26.4 41.7
SNAC (24kHz) 0.98 2.25 1.48 2.23 0.914 - - - 35.4 16.7
WavTokenizer (Small, F=75F=75Hz) 0.90 2.45 1.63 2.47 0.925 - - - 37.2 15.5
Stable-Codec (2 residual FSQ) 0.70 4.94 1.73 2.16 0.917 - - - NC 14.0
Acoustic
(Semantic Distillation)
SpeechTokenizer 4.00 1.47 1.20 2.60 0.930 - - - 19.3 57.3
X-Codec (HuBERT LibriSpeech) 4.00 1.27 1.49 2.82 0.905 - - - 9.8 69.6
Mimi 1.10 2.44 1.68 2.27 0.917 - - - 26.8 50.9
LLM-Codec 0.85 6.25 1.86 1.82 0.879 - - - NC 16.2
[…]
Acoustic
FACodec (with detail codes) 4.80 1.37 1.02 2.91 0.954 1.57 0.773 0.583 14.6 51.1
LSCodec (F=50F=50Hz) 0.45 3.33 2.42 1.77 0.688 4.04 0.852 0.697 25.3 49.8
Semantic (SSL) vq-wav2vec (k-means) 1.80 2.81 2.73 1.49 0.795 3.27 0.857 0.718 16.9 58.7
wav2vec 2.0 (Large, inner quantizer) 0.90 3.24 2.92 1.52 0.680 4.40 0.814 0.759 22.0 51.9
wav2vec 2.0 (Large, L14+km2048) 0.55 2.51 9.57 1.20 0.630 2.81 0.880 0.492 5.8 69.5
HuBERT (Large, L24+km2048) 0.55 1.86 15.65 1.17 0.625 1.97 0.876 0.375 6.1 67.2
WavLM (Large, L24+km2048) 0.55 1.67 17.94 1.16 0.621 1.92 0.872 0.374 6.1 74.2
ContentVec (L12+km2048) 0.55 2.09 18.88 1.15 0.613 2.21 0.869 0.348 5.5 72.0
Semantic (Supervised) $\mathcal{S}^{3}$ Tokenizer (F=50F=50Hz) 0.60 2.12 4.25 1.37 0.673 2.52 0.868 0.687 17.5 67.2
To enable a fair comparison between acoustic and semantic tokens from a reconstruction perspective, we train a CTX-vec2wavα vocoder [guo2024lscodec] for different semantic tokens on LibriTTS [libritts]. This vocoder supplements the insufficient speaker timbre information in semantic tokens using continuous WavLM features extracted from the reference prompts. This approach enables semantic tokens to perform voice conversion (VC) by switching reference prompts conveniently. The training details follow [du2024unicats]. We compute several metrics for reconstruction ability:
@@
We use LibriTTS testset-B  [du2024unicats] as the test set for evaluations. It contains 500 utterances from unseen speakers that sum up to about 1 hour. We use the original utterance to provide timbre information when necessary, i.e. for TiCodec, FACodec, LSCodec and all semantic tokens. All evaluation metrics are computed on 16kHz waveform for a fair comparison, and reconstructed waveforms with higher sampling rates are downsampled before evaluation.
@@
The results are shown in Table I. It is evident that acoustic tokens designed only for reconstruction can achieve decent speech quality, but still far from the state-of-the-art spectrogram-based vocoders because of higher compression rates. Retaining good speech intelligibility (i.e. low WER) becomes particularly challenging when the frame rate is low. Acoustic tokens with semantic distillation can also achieve strong reconstruction quality. Explicitly disentangled acoustic tokens may sacrifice some reconstruction performance metrics when the bitrate is extremely low. Semantic tokens generally struggle to achieve the same level of acoustic reconstruction as acoustic tokens, as evidenced by lower GPE, PESQ, and STOI scores. Notably, most semantic tokens included exhibit significant information loss in prosody as reflected by their GPE scores. However, their WER scores remain comparable to acoustic tokens, despite having much lower bitrates. This highlights the property that semantic tokens primarily retain content-related information rather than acoustic details.
VI-D Voice Conversion Analysis